OpenStudy (anonymous):

Find two numbers with a sum of -6 and a difference of 34.

OpenStudy (anonymous):

X+y = -6 x-y = 34 x= -6-y (-6-y)-y=34 -2y=40 y=-20 x+(-20)=-6 x=14 y=-20 and x=14 :-)